# Emergent predictability in microbial ecosystems

### v2

@@ Abstract
Microbial ecosystems carry out essential functions for global climate, human health, and industry. These complex communities exhibit a surprising amount of functionally relevant diversity at all levels of taxonomic resolution, presenting a significant challenge for most modeling frameworks. A long-standing hope of theoretical ecology is that some patterns might persist despite community complexity – or perhaps even emerge because of it. A deeper understanding of such “emergent simplicity” could enable new approaches for predicting the behaviors of the complex ecosystems in nature. However, most examples described so far afford limited predictive power, as they focused on reproducibility rather than prediction. Here, we propose an information-theoretic framework for defining, nuancing and quantifying emergent simplicity in empirical data based on the ability of simple models to predict community-level functional properties. Applying this framework to two published datasets, we demonstrate that the majority of properties measured across both experiments exhibit robust evidence of emergent predictability: surprisingly, as community richness increases, simple compositional descriptions become more predictive. We show that this behavior is not typical within the standard modeling frameworks of theoretical ecology, and argue that improving our ability to predict and control natural microbial communities will require a shift of focus: away from complexity of ecosystems, and towards prediction complexity of properties of ecosystems.
@@
Microbial communities perform key functions in a wide range of contexts, including human health, agriculture, and the global climate. A key goal of microbial ecology has thus been to predict community function from its composition. Recent advances in high-throughput sequencing technologies provide a high taxonomic resolution of the composition of microbial communities, revealing that these communities consistently exhibit high richness (number of coexisting taxa), a surprising diversity at all levels of resolution, high functional redundancy, and a lack of reproducibility in even similar conditions. These characteristics pose significant challenges for making quantitative predictions, which are necessary for applications in industry and medicine.
@@
A long-standing hope of theoretical ecology is that, despite this complexity, some patterns might nevertheless be predictable [1, 2, 3]. The search for emergent regularities in complex ecosystems is as old as ecology itself. For example, classical work uncovered reproducible trophic structures across taxonomically divergent communities [4] or a consistent ratio of predator species to prey species [5, 6]. More recently, moving beyond the patterns persisting despite high ecosystem diversity, it was proposed that some regularities could perhaps arise because of it [7, 8, 9], coining the term “emergent simplicity” [7, 10, 11, 12, 13]. This notion is at the foundation of the interface between statistical physics and the so-called large-$N$ ecology, explaining species abundance distributions and other patterns as some parameter becomes large [14, 15, 16, 17, 1, 8, 18, 19].
Figure 1: A quantitative framework for interrogating emergent simplicity in data. A: A general experimental protocol with community richness (number of strains) as a control parameter. A fixed library of strains is subsampled to assemble communities of low richness (top) or high richness (bottom). Community composition (strain abundances, $n_{i}$) and some property of interest $Y$ are measured in each community. B: Our approach will be to train models predicting $Y$ based on combined abundances of groups of taxa. Each choice of a coarsening $\Psi$ defines a regression model (coefficients $\{\beta_{i}^{\Psi}\}$), whose performance we will evaluate. C: For each coarsening $\Psi$, we can evaluate the description complexity $I(\Psi)$ (the extent of coarsening; see text) and the out-of-sample prediction error, quantified as the fraction of variance unexplained (FVU) by the corresponding regression model. The cartoon illustrates the expected general shape of the scatter plot in these axes, evaluated for all possible coarsenings $\Psi$ (points). Our focus is on the shape of the Pareto front (thick line), characterizing the best predictive power achievable by models of a given complexity. As community richness increases, the phenomenon of emergent simplicity is identified by the behavior of the Pareto fronts, and can take one of two distinct forms: D: emergent coarse-grainability (horizontal arrow), or E: emergent predictability (vertical arrow). We stress that our normalization—focusing on the fraction of variance unexplained, equivalent to normalizing all observables to unit variance—ensures that both forms of emergent simplicity defined here are distinct from the simplest law-of-large-numbers intuition (a.k.a. central limit theorem or “self-averaging”); see text.
@@
The idea of emergent simplicity is very appealing, suggesting that functional predictions could perhaps be made directly in the complex natural context [20]. However, so far this hope remains out of reach. Most reported instances of “emergent simplicity” fall under the umbrella of functional convergence or self-averaging, focusing on reproducibility rather than prediction [4, 5, 6, 21, 7, 10, 22, 23, 11, 12, 13]. Unfortunately, such reproducibility is often approximate or limited to simplified conditions [24], and even when it holds, offers limited predictive power. As an illustration, consider the statement: “if the soil ecosystem is diverse enough, the rate of process $P$ will be $\approx\!\!X$ per gram of biomass”. This blanket expectation would likely be somewhat naïve, but even if it were true, it would leave key questions unanswered – such as how the process rate would be affected by a perturbation, or what intervention would help increase it. As a result, observing emergent regularities of this kind has generally not led to improved functional predictions.
@@
Understanding and harnessing emergent simplicity could transform our ability to predict and control complex natural communities. However, to realize this potential we require a framework focused on prediction rather than reproducibility. Here, we introduce an information-theoretic framework to explicitly quantify the predictive power of simple models in ecological data. This allows us to convert the aspiration of emergent simplicity into a testable quantitative hypothesis. Applying our framework to two published datasets on synthetic ecosystems, we show that, remarkably, the predictive power of simple models indeed improves with community richness, this improvement strongly supported for 4 out of 5 properties measured across both experiments. These results provide support for emergent simplicity as a valid path towards predictions, one for which community richness is an asset rather than a nuisance.
@@ I.1 A model experimental setup
For concreteness, consider the experimental setup depicted in Fig. 1A. Let $\mathcal{L}$ be a fixed library of $S$ strains, indexed by $i$. Subsets of these strains are sampled from this library to inoculate $N$ communities, indexed by $\mu$ and defined by a binary matrix $b_{i\mu}$ identifying whether strain $i$ is included in community $\mu$. We will assume that community richness $R_{\mu}\equiv\sum_{i}b_{i\mu}$ is kept consistent across the dataset; this richness will become our key control parameter for comparisons (a low-richness dataset vs. a high-richness dataset). Once the communities are assembled, we measure their microscopic composition (abundance of each strain $n_{i\mu}$) as well as some community-level property of interest $Y_{\mu}$ that we seek to predict (e.g., production of a metabolite). We will not assume that the communities reach any notion of a steady state, merely that some assembly protocol is followed consistently in each trial.
The simplest, and most familiar, notion of emergent simplicity is self-averaging. As community richness increases, one might observe the distribution of $Y_{\mu}$ become increasingly tight. Much previous work suggests that certain properties of high-richness communities may indeed exhibit such “functional convergence.” [4, 5, 6, 21, 7, 10, 22, 23, 11, 12, 13, 8] However, for the purposes of this work, what we care about is not the total variance of $Y_{\mu}$, but our ability to explain it from community composition. Our goal is to quantify the predictive power of simple models in ecological datasets, and then use empirical data to ask how this predictive power changes with community richness.
@@ I.2 The candidate ‘simple models’
Our approach is motivated by an empirical example, namely wastewater treatment bioreactors that use microbes to digest organic waste into methane. These complex communities assemble semi-naturally and can harbor over 1000 taxa [25]. The rate of digestion (the amount of methane produced) can vary widely [26, 27], but, remarkably, can be explained by simple models coarse-graining this vast diversity into just 4 functional classes (acidogens, propionate-degrading acetogens, butyrate-degrading acetogens, and methanogens [28]).
@@
Motivated by this example, throughout this work, our candidate simple models will be indexed by the choice of a coarsening $\Psi$, which maps individual strains into a smaller number of classes: $[1\dots S]\xrightarrow{\Psi}[1\dots K^{\Psi}]$ with $K^{\Psi}\leq S$ (Fig. 1A). Some natural choices of coarsening $\Psi$ might be to group strains based on taxonomy or key traits, but we will allow any choice of $\Psi$ as valid.
@@
For any coarsening $\Psi$, we will assess the prediction error of the best model from a specified model class $\mathcal{M}$, taking as input the combined abundances $\tilde{n}^{\Psi}_{j\mu}=\sum_{\Psi(i)=j}n_{i\mu}$. This approach echoes other recent attempts to relate community-level properties to coarse-grained compositional features [29, 30]. The model class $\mathcal{M}$ could be, for instance, the class of Generalized Lotka-Volterra models, or some general-purpose ansatz (e.g., training a neural network). In this work, we will pick a particularly simple class of models, namely linear regressions $Y_{\mu}=\beta^{\Psi}_{0}+\sum_{i}\beta^{\Psi}_{i}\tilde{n}^{\Psi}_{i\mu}+\epsilon_{\mu}$ (where $\epsilon_{\mu}$ is the residual noise). This choice is motivated by two considerations. First, we will want to apply our framework to data, and as we will see, the limitations of existing datasets make it preferable to pick a model class trainable on fewest data points. Second, it was recently shown that simple regressions can in fact be surprisingly predictive of ecosystem function [31], motivating us to proceed.
In summary, the candidate “simple models” we will consider are indexed by the choice of a coarsening $\Psi$. Each $\Psi$ defines a set of coarsened variables ($\tilde{n}^{\Psi}_{i\mu}$), which allow us to train an instance of a linear regression model (coefficients $\beta^{\Psi}_{i}$), whose prediction error we will evaluate.
@@
It is worth noting that a common approach to vary modeling complexity is to keep the variables fixed (e.g. “foxes”, “rabbits”), but change the model class (e.g, a non-interacting model; adding quadratic Lotka-Volterra interactions; then higher-order terms; etc.). In contrast, the defining feature of our approach is to fix the model class, and change the variables (vary the extent of taxon grouping into putative functional classes). Our next step is to define quantitative metrics for the performance and complexity of each model thus constructed.
@@ I.3 Quantifying the performance of simple models
For each candidate “simple model” (coefficients $\beta^{\Psi}_{i}$) indexed by the coarsening $\Psi$, we will evaluate two features. First, we will compute its out-of-sample prediction error. Since we explicitly want to go beyond simple self-averaging, we will normalize away the overall width of the distribution of $Y$, and report the fraction of unexplained variance (FVU), defined as the mean square error of prediction divided by the variance of $Y$, and confined between 0 and 1.
Second, we will evaluate the description complexity $I(\Psi)$, which we define as the mutual information between the label of the coarse-grained class and the identity of the microscopic strain:
$I(\Psi)=-\sum_{j=1}^{K^{\Psi}}\frac{S^{\Psi}_{j}}{S}\log\frac{S^{\Psi}_{j}}{S}.$
Here $S_{j}^{\Psi}$ is the number of strains that $\Psi$ assigns to class $j$; $K^{\Psi}$ is the total number of classes; and $S$ is the total number of strains in the library. Our definition of $I(\Psi)$ implements the intuition that description complexity should grow with the number of classes it distinguishes, but generalizes it to groups of unequal size (for additional discussion, see SI section “Quantifying description complexity”). This metric $I(\Psi)$ reaches its maximum for the fully microscopic description, when each strain is its own class: $K^{\Psi}=S$ and $I(\Psi)=\log S$.
Established results of statistical learning theory indicate that if we could evaluate and plot the fraction of unexplained variance against the description complexity $I(\Psi)$ for all possible coarsenings $\Psi$, the result would be expected to look like Fig. 1C. Each point in the point cloud corresponds to one choice of $\Psi$, and describes the predictive power of one candidate simple model.
@@
The left-most point $\Psi_{z}$ is the ‘zero-information’ description that groups all strains together. If no compositional information is available (only the total biomass), we expect the fraction of unexplained variance—the ordinate of $\Psi_{z}$—to be close to 1. (If $\Psi_{z}$ is notably below 1, i.e., the chosen property obviously scales with biomass, it might be more meaningful to seek to predict its value per unit biomass; see SI section “Normalization of predictor and response variables”.) As the information content of a description increases, better predictions become possible. Models that provide the lowest prediction error for a given complexity constitute the Pareto front (thick line). The Pareto front continues to trend downwards until descriptions become too complex to be fit on limited data (overfitting), and out-of-sample prediction error starts increasing (Fig. 1C). From now on, our focus will be on the properties of this Pareto front.
I.4 Distinguishing emergent coarse-grainability and emergent predictability
By construction, the point cloud in Fig. 1C, and thus the shape of the Pareto front, explicitly depend on the dataset on which the predictive power is being evaluated. Our goal is to compare the Pareto front constructed for a dataset consisting of high-richness examples vs. low-richness examples (Fig. 1A).
We distinguish two behaviors of interest (Fig. 1D). As community richness is increased, we might observe the Pareto front shifting to the left. This would mean that the same prediction accuracy can be achieved by increasingly coarse descriptions (‘emergent coarse-grainability’). Separately, we might observe the Pareto front shifting down. This would mean that models of comparable complexity become increasingly predictive (‘emergent predictability’).
Both phenomena could be collated under the umbrella term ‘emergent simplicity,’ but are clearly distinct. A given system could exhibit only one of these, or both, or neither. Having built a theoretical framework to distinguish them, we now turn to asking whether these behaviors are expected (based on models), and whether they are observed (based on empirical data).

### v3

@@ Abstract
Microbial ecosystems exhibit a surprising amount of functionally relevant diversity at all levels of taxonomic resolution, presenting a significant challenge for most modeling frameworks. A long-standing hope of theoretical ecology is that some patterns might persist despite community complexity – or perhaps even emerge because of it. A deeper understanding of such “emergent simplicity” could enable new approaches for predicting the behaviors of the complex ecosystems in nature. However, the concept remains partly intuitive with no consistent definition, and most empirical examples described so far afford limited predictive power. Here, we propose an information-theoretic framework for defining and quantifying emergent simplicity in empirical data based on the ability of coarsened descriptions to predict community-level functional properties. Applying this framework to two published datasets, we demonstrate that all five properties measured across both experiments exhibit robust evidence of what we define as “emergent predictability”: surprisingly, as community richness increases, simple compositional descriptions become more predictive. We show that standard theoretical models of high-diversity ecosystems fail to recapitulate this behavior. This is in contrast to simple self-averaging, which is well-understood and generic across models. We propose that, counterintuitively, emergent predictability arises when physiological or environmental feedbacks oppose statistical self-averaging along some axes of community variation. As a result, these axes of variation become increasingly predictive of community function at high richness. We demonstrate this mechanism in a minimal model, and argue that explaining and leveraging emergent predictability will require integrating large-$N$ theoretical models with a minimal notion of physiology, which the dominant modeling frameworks currently omit.
@@
Microbial communities perform key functions in a wide range of contexts, including human health, agriculture, and the global climate. Predicting community function from its composition is a key goal of microbial ecology. Recent advances in high-throughput sequencing allowed compositional information to be obtained in high resolution, and revealed that natural communities consistently exhibit high richness (number of coexisting taxa), a surprising diversity at all levels of resolution, high functional redundancy, and a lack of reproducibility in even similar conditions. These characteristics pose significant challenges for making quantitative predictions, which are necessary for applications in industry and medicine.
@@
A long-standing hope of theoretical ecology has been that, despite this complexity, some patterns might nevertheless be predictable [37, 72, 71]. The search for emergent regularities in complex ecosystems is as old as ecology itself. For example, classical work uncovered reproducible trophic structures across taxonomically divergent communities [25] or a consistent ratio of predator species to prey species [12, 8]. More recently, moving beyond the patterns persisting despite high ecosystem diversity, it was proposed that some regularities could perhaps arise because of it [22, 20, 3], coining the term “emergent simplicity” [22, 18, 42, 48, 10]. This notion is at the foundation of the interface between statistical physics and the so-called large-$N$ ecology, explaining species abundance distributions and other patterns as some parameter (e.g., the number of species) becomes large [1, 39, 15, 38, 37, 20, 51, 52].
@@
The idea of emergent simplicity is very appealing, suggesting that functional predictions could perhaps be made directly in the complex natural context [39, 6]. However, so far this hope remained out of reach. In the absence of a consistent definition, it is unclear precisely what forms of “emergent simplicity” should be seen as surprising or useful. Many reported instances fall under the umbrella of functional convergence, focusing on reproducibility rather than prediction [25, 12, 8, 16, 22, 18, 55, 56, 42, 48, 10]. Such reproducibility can be difficult to distinguish from simple self-averaging or the central limit theorem [41, 27] and offers limited predictive power. As an illustration, consider the statement: “if the soil ecosystem is diverse enough, the rate of process $P$ will be $\approx\!\!X$ per gram of biomass”. This blanket expectation would likely be somewhat naïve, as empirical observations of reproducibility are often approximate or limited to simplified conditions [66]. But even if it were true, it would leave key questions unanswered, such as how the process rate would be affected by a perturbation, or what intervention would help increase it. As a result, observing emergent regularities of this kind has generally not led to improved functional predictions.
@@
Understanding and harnessing emergent simplicity could transform our ability to predict and control complex natural communities. However, to realize this potential we need to make this concept well-defined, measurable, and focused on prediction rather than reproducibility. Here, we introduce a theoretical framework to explicitly quantify the predictive power of coarsened compositional descriptions in ecological data. This allows us to convert the aspiration of emergent simplicity into a testable quantitative hypothesis. Applying our framework to two published datasets on lab-assembled ecosystems, we show that, remarkably, for all 5 properties measured across both experiments, the predictive power of coarse descriptions indeed improves with community richness. We further show that standard theoretical models of high-diversity ecosystems fail to recapitulate this behavior. This is in contrast to simple self-averaging, which is well-understood and generic across models. We propose that, counterintuitively, emergent predictability arises when physiological or environmental feedbacks oppose statistical self-averaging along some axes of community variation. As a result, these axes of variation can become increasingly predictive of community function at high richness. We demonstrate this mechanism in a minimal model, and argue that explaining and leveraging emergent predictability will require integrating large-$N$ theoretical models with a minimal notion of physiology, which the dominant modeling frameworks currently omit.
@@ I.1 A model experimental setup
For concreteness, consider the experimental setup depicted in Fig. 1A. Let $\mathcal{L}$ be a fixed library of $S$ strains, indexed by $i$. Subsets of these strains are sampled from this library to inoculate $N$ communities, indexed by $\mu$ and defined by a binary matrix $b_{i\mu}$ identifying whether strain $i$ is included in community $\mu$. We will assume that community richness $R_{\mu}\equiv\sum_{i}b_{i\mu}$ is kept consistent across the dataset; this richness will become our key control parameter for comparisons (a low-richness dataset vs. a high-richness dataset). Once the communities are assembled, we measure their microscopic composition (abundance of each strain $n_{i\mu}$) as well as some community-level property of interest $Y_{\mu}$ that we seek to predict (e.g., production of a metabolite). We will not assume that the communities reach any notion of a steady state, merely that some assembly protocol is followed consistently in each trial. Our goal is to quantify the predictive power of simple (coarsened) descriptions in ecological datasets of this kind, and then use empirical data to ask how this predictive power changes with community richness.
Figure 1: A quantitative framework for interrogating emergent simplicity in data. A: A general experimental protocol with community richness (number of strains) as a control parameter. A fixed library of strains is subsampled to assemble communities of low richness (top) or high richness (bottom). Community composition (strain abundances, $n_{i}$) and some property of interest $Y$ are measured in each community. B: Our approach will be to train models predicting $Y$ based on combined abundances of groups of taxa. Each choice of a coarsening $\Psi$ defines a regression model (coefficients $\{\beta_{i}^{\Psi}\}$), whose performance we will evaluate. C: For each coarsening $\Psi$, we can evaluate the description complexity $I(\Psi)$ (the extent of coarsening; see text) and the out-of-sample prediction error, quantified as the variance unexplained by the corresponding regression model. The cartoon illustrates the expected general shape of the scatter plot in these axes, evaluated for all possible coarsenings $\Psi$ (points). Our focus is on the shape of the Pareto front (thick line), characterizing the best predictive power achievable by models of a given complexity. The behavior of the Pareto fronts defines three distinct (but not mutually exclusive) forms of “emergent simplicity”: D: self-averaging, when higher community richness leads to a tighter distribution of property $Y$; E: coarse-grainability, when coarsened descriptions are surprisingly predictive (compared to a null model), at fixed richness; or F: emergent predictability, when the predictive power of coarsened descriptions improves with ecosystem richness.
@@ I.2 The candidate ‘simple models’
Our approach is motivated by an empirical example, namely wastewater treatment bioreactors that use microbes to digest organic waste into methane. These complex communities assemble semi-naturally and can harbor over 1000 taxa [31]. The rate of digestion (the amount of methane produced) can vary widely [34, 49], but, remarkably, can be largely explained by simple models coarse-graining this vast diversity into just 4 functional classes (acidogens, propionate-degrading acetogens, butyrate-degrading acetogens, and methanogens [7]).
@@
Motivated by this example, throughout this work, our candidate simple models will be indexed by the choice of a coarsening $\Psi$, which maps individual strains into a smaller number of classes: $1\dots S]\xrightarrow{\Psi}[1\dots K^{\Psi}]$ with $K^{\Psi}\leq S$ (Fig. 1B). Some natural choices of coarsening $\Psi$ might be to group strains based on their taxonomy or key traits, but we will allow any choice of $\Psi$ as valid, and define description complexity $I(\Psi)$ as the mutual information between the label of the coarse-grained class and the identity of the microscopic strain:
$I(\Psi)=-\sum_{j=1}^{K^{\Psi}}\frac{S^{\Psi}_{j}}{S}\log\frac{S^{\Psi}_{j}}{S}.$
Here $S_{j}^{\Psi}$ is the number of strains that $\Psi$ assigns to class $j$; $K^{\Psi}$ is the total number of classes; and $S$ is the total number of strains in the library. Our definition of $I(\Psi)$ implements the intuition that description complexity should grow with the number of classes it distinguishes, but generalizes it to groups of unequal size (for additional discussion, see Supplementary Text, section B.1). This metric $I(\Psi)$ reaches its maximum for the fully microscopic description, when each strain is its own class: $K^{\Psi}=S$ and $I(\Psi)=\log S$.
@@
For any coarsening $\Psi$, we will assess the prediction error of the best model from a specified model class $\mathcal{M}$, taking as input the combined abundances $\tilde{n}^{\Psi}_{j\mu}=\sum_{\Psi(i)=j}n_{i\mu}$ and predicting $Y$. This approach echoes other recent attempts to relate community-level properties to coarse-grained compositional features [65, 47]. The model class $\mathcal{M}$ could be mechanistically motivated (for instance, the class of Generalized Lotka-Volterra models predicting a dynamical property like future abundance of a focal species), or some general-purpose ansatz (e.g., a neural network). In this work, we will pick a particularly simple class of models, namely linear regressions $Y_{\mu}=\beta^{\Psi}_{0}+\sum_{i}\beta^{\Psi}_{i}\tilde{n}^{\Psi}_{i\mu}+\epsilon_{\mu}$ (where $\epsilon_{\mu}$ is the residual noise). This choice is motivated by two considerations. First, we will want to apply our framework to data, and as we will see, the limitations of existing datasets make it preferable to pick a model class trainable on fewest data points. Second, it was recently shown that simple regressions can in fact be surprisingly predictive of ecosystem function [67, 3], motivating us to proceed.
@@
In summary, the candidate “simple models” we will consider are indexed by the choice of a coarsening $\Psi$. Each $\Psi$ defines a set of coarsened variables ($\{\tilde{n}^{\Psi}_{i\mu}\}$), which allow us to train an instance of a linear regression model (coefficients $\beta^{\Psi}_{i}$), whose prediction error we will evaluate. It is worth noting that a common approach to vary modeling complexity is to keep the variables fixed (e.g. “foxes”, “rabbits”), but change the model class (e.g, a non-interacting model; adding quadratic Lotka-Volterra interactions; then higher-order terms; etc.). In contrast, the defining feature of our approach is to fix the model class, and change the variables (vary the extent of taxon grouping into putative functional classes).
@@ I.3 Quantifying the performance of simple models
For each candidate “simple model” (coefficients $\beta^{\Psi}_{i}$) indexed by the coarsening $\Psi$, we can quantify its prediction error, computed on held-out data. The well-established bias-variance tradeoff from statistical learning theory indicates that if we could evaluate and plot the prediction error (the variance left unexplained) against the description complexity $I(\Psi)$ for all possible coarsenings $\Psi$, the result would be expected to look like Fig. 1C. Each point in the point cloud corresponds to one choice of $\Psi$, and describes the predictive power of one candidate simple model. The left-most point $\Psi_{B}$ is the least predictive, coarsest description that groups all strains together ($I(\Psi_{B})=0$). If no compositional information is available—only the total biomass—the unexplained variance of $Y$ is close to its total variance. (For properties that scale with biomass, it might be more meaningful to seek to predict their value per unit biomass; see Supplementary Text, section B.2.) As the information content of a description increases, better predictions become possible. Models that provide the lowest prediction error for a given complexity constitute the Pareto front (thick line). The Pareto front continues to trend downwards until descriptions become too complex to be fit on limited data (overfitting), and out-of-sample prediction error starts increasing (Fig. 1C). From now on, our focus will be on the behavior of this Pareto front.
I.4 Distinguishing self-averaging, coarse-grainability and emergent predictability
The intuitive notion of “emergent simplicity” can be codified in three distinct ways. These are important to distinguish, since as we will see, they have very different properties.
The simplest, and most familiar, version of emergent simplicity is self-averaging. As community richness increases, one might observe the distribution of $Y_{\mu}$ become increasingly tight. Much previous work indicates that properties of high-richness communities may indeed exhibit such “functional convergence.” [25, 12, 8, 16, 22, 18, 55, 56, 42, 48, 10, 20] We can think of this width as the prediction error of a particularly trivial model which predicts a constant value, such that the unexplained variance is synonymous with the total variance.
In practice, the predictions that interest us are rarely of this “zero-information” kind. Instead, we usually seek to predict properties of a community based on its composition. To define notions of simplicity that go beyond simple self-averaging, it is convenient to normalize away the overall width of the distribution of $Y$, and think of the fraction of unexplained variance (FVU), defined as the mean square error of prediction divided by the variance of $Y$.
In these normalized units, the Pareto front descends from approximately 1 (no compositional information; worst prediction) to the best prediction attainable with this model class (in the limit of infinite data—neglecting overfitting—this would correspond to the full microscopic description). How quickly it descends encodes the notion of “coarse-grainability”. For some properties, we may find that a surprisingly coarse description already enables good predictions; this form of simplicity has also been widely reported [21, 36, 57, 64]. To rigorously define this concept, it is essential to quantify the word “surprisingly,” defining a null expectation to which the shape of this curve should be compared. Our framework will enable us to do so, allowing us to examine “coarse-grainability” of ecological properties in empirical data.
Finally, we can ask what happens to the Pareto front when we change ecosystem richness, and compare the fronts constructed for a dataset restricted to only high-richness examples vs. only low-richness examples, for the same property $Y$ (Fig. 1A). A consistent shift of the Pareto front would indicate that, as richness is increased, models of comparable complexity become increasingly predictive (Fig. 1F, arrow “a”), and/or that same prediction accuracy can be achieved by increasingly coarse descriptions (Fig. 1F, arrow “b”). While the two directions of shift have different verbal interpretations, in general they cannot be formally distinguished, and we will instead quantify “emergent predictability” by measuring the area contained between the two Pareto fronts (see Methods).
The rest of the paper can be summarized in two claims: (1) Self-averaging, coarse-grainability and emergent predictability are all observed in empirical data. However, (2) the three phenomena differ dramatically in their origins, properties and the attention they deserve. Today, the broad usage of the phrase “emergent simplicity” came to encompass any observation of something simple in a complex community context. In this sense, all three phenomena just defined could be collated under this umbrella term. However, as we will show, self-averaging is generic and ubiquitous; on its own, it is neither surprising nor special. Coarse-grainability does not come “for free” (from diversity alone), but can be a straightforward consequence of the fact that some species are more similar than others. While missing from some widely used models that omit such structure, its empirical observation should also not be seen as surprising, unless the observation comes with a quantitative argument that the observed coarse-grainability is stronger than expected. In contrast, emergent predictability is missing from both random and structured theoretical models – and yet, as we will see, it is strongly supported by data. This is the phenomenon that we propose should be seen as “emergent simplicity” sensu stricto. Its empirical observation is surprising and demands a theoretical explanation, and we will present evidence supporting a candidate mechanism for its emergence.